OpenStudy (anonymous):

what is the domain for f(x)=x(7-2x)^2

hartnn (hartnn):

when there is denominator, all values of x when denominator =0 are excluded from domain when there is radical \((\sqrt {stuff})\), domain will have stuff >=0 when the function is polynomial, the domain is all real numbers.

hartnn (hartnn):

so, now can you tell me whats the domain of your function ?

OpenStudy (anonymous):

i thought it was all real numbers, but when i plugged it in to my hw site it said it was wrong

hartnn (hartnn):

are tou sure the question looked like \(\Large f(x)=x (7-2x)^2\) ?? if yes, then the domain IS all real numbers, for sure.

OpenStudy (anonymous):

yeah that is the equation but it won't accept all real numbers as the correct answer.

hartnn (hartnn):

can you attach a screenshot ? maybe they want answer in another form...

OpenStudy (anonymous):

it asked for it in interval notation, it was originally a word problem which i solved for and got the equation that i sent you, it said that part was correct online.

hartnn (hartnn):

interval notation, \(\Large (-\infty,\infty)\) did u put this ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

that domain in interval notation is surely correct for the function you mentioned. i think there's some error in your system....ask your teacher about it.

OpenStudy (anonymous):

i will, thank you for your help.

hartnn (hartnn):

you're welcome ^_^ and since you are new here, if you have any doubts browsing this site, you can ask me :) \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)